# SpecGen: Automated Generation of Formal Program Specifications via Large Language Models

### v4

@@ I Introduction
  • A novel approach for formal program specification generation leveraging the Large Language Models to generate accurate and comprehensive specifications to describe program behaviors. Benefiting from the code comprehension ability of LLMs, our approach is capable of generating specifications with high quality, overcoming the limitations of existing methods in generating simplistic and basic specifications.
@@
  • A comprehensive evaluation from different aspects to evaluate the effectiveness of our approach. We compare SpecGen against Purely LLM-based approaches and representative non-LLM approaches. SpecGen succeeds in 279 out of the 385 programs, significantly outperforming the baseline approaches.
@@ II-A Specification Generation and Verification
Program specifications encompass precise statements that describe the intended or actual behaviors of a particular program, either in its entirety or in distinct parts. In this work, we focus on generating specifications for the actual behaviors of input programs. A large proportion of program specifications are expressed in formal languages, such as mathematical expressions to describe the constraints on the behaviors of a program. There are different kinds of specifications such as pre-conditions, which establish constraints on function parameters, ensuring proper execution of the function, post-conditions, which delineate the properties of a set of variables that persist after a function is executed, and loop invariants, which represent a specialized form of specification, detailing properties that consistently hold before executing the loop body. For different programming languages, the specifications may have different implementation forms. For example, in Java, the specifications can be expressed in Java Modeling Language (i.e., JML) [22] where requires statements denote the pre-conditions of a function, ensures statements represent the post-conditions of a function, and maintaining statements are the loop invariants in the program.
@@ II-B Motivation
Recently, large language models (i.e., LLMs) [27, 28] have exhibited powerful capacities in coding [19, 18]. The emergence of these models may greatly compensate for the limitations of traditional software analysis tools in code understanding. A significant amount of work attempts to leverage large language models in software engineering [29, 30, 31, 32, 33] and we have witnessed substantial progress brought about by the introduction of LLMs. Inspired by these works, in this work, we aim to leverage the large language models in the automated generation of formal program specifications to address the limitations of conventional template-based approaches. From this perspective, we innovate our approach SpecGen, which generated three parts of specifications for the example presented in the right part of Figure 1. The first part is to describe the method TwoSum() as a whole, specifying its pre-conditions and post-conditions. The specifications in lines a and b claim that the input array must not be null before and after the method is executed. The post-condition in line c specifies that the target value equals the sum of the two elements corresponding to the indexes stored in the returned array. The post-condition at line d specifies that there does not exist such a pair of elements that satisfies the constraint when the length of the returned array is zero. The functionality of method TwoSum() can be fully articulated by these generated specifications. Furthermore, the loop invariants in the second and third parts specify corresponding constraints that must be met within a certain range in the array. These generated specifications by SpecGen comprehensively describe the semantics of this function and their correctness is verified by the verifier.
@@ III-A Overview
The overview of SpecGen is presented in Figure 3, which consists of two components i.e., conversation-driven specification generation and mutation-based specification generation. For the former, SpecGen is designed to communicate with the large language model to query the output in a conversational manner. In particular, a prompt is constructed with some few-shot examples for the initial query. The verification failure information provided by the verifier is further used as the prompt for the next round of conversation if the model-generated results are incorrect. The conversation will be repeated iteratively until the generated specifications successfully pass the verifier or a maximum number of iterations is reached. For the latter, which aims at generating the specifications of a program that the large language model fails to generate. Four kinds of mutation operators are adopted to mutate the specification that failed verification by the verifier and obtain all potential variants. A heuristic selector is further designed to efficiently choose a set of mutated variants that is most likely to pass verification.
@@ III-C Mutation-based Specification Generation
In the conversation-driven generation process, some few-shot examples are provided in the initial prompt to start the query with the LLM. To further stimulate the potential of the LLM, multi-turn conversation continually guides the LLM to approach the accurate specification more closely. Yet, they still struggle to generate fully correct specifications for some complex programs. The reasons are two-fold. On one hand, in comparison to code generation [36, 37, 38], specification generation poses greater challenges, even for LLMs, because there are currently limited corpora related to program specifications for the model to learn from. On the other hand, although the verification failure information provided by the verifier can assist LLMs in providing higher-quality responses to some extent, as the error messages are highly abstract and generalized, LLMs still struggle to accurately understand the semantic information within error messages for complex programs, leading to incorrect specifications. While LLMs may not accurately generate specifications for complex programs, we found that the generated results are already close enough to the oracle, inspiring us to design the mutation-based generation component.
@@ III-C1 Template Specification Mutation
As shown in Table I, we define four kinds of mutation operators including predicative, logical, comparative, and arithmetic. Each type of mutation corresponds to one type of operator supported by JML. LLMs perform well in capturing the overall semantics of a program, but they often make mistakes in grasping the fine-grained relationships between variables, resulting in incorrect operators used to describe the relationships between variables, which is why our mutation design is centered around the operators. When applying the mutation operator to a specification, it substitutes the operators of the corresponding type in the specification with another operator of the same type. For example, after applying a predicative mutation, a $\exists$ predicate within a specification may be substituted with $\forall$. Note that the mutation for a certain type of operator does not necessarily create only one mutated candidate. For example, the expression a <= b may be mutated to a < b or a - 1 <= b. If multiple mutations can be applied to a specification at the same time, we try to exhaust each combination of different types of mutations to get all potential variants. Since the set of all potential variants of a certain template is determined, the exhaustive searching process is deterministic. For instance, the expression x < n + 1 can be mutated to x <= n + 1 from the comparative type, x < n - 1 from the arithmetic type, or x <= n - 1 by combining them.
@@ III-C2 Mutated Specification Selection
Typically, for a program, Houdini [7] usually verifies all generated specifications at one time. However, a similar practice cannot be applied in SpecGen as we try to collect all potential combinations of mutations for a template specification. The set of the obtained specifications for verification is considerably large, which poses a much greater burden for the verifier within a single verification process. To address this challenge, we innovate a heuristic selection strategy to improve the stability and efficiency of verification.
@@ IV-B Dataset
As Daikon [9] requires a set of test suites instrumented into the source code to execute the code, we manually write these test suites for it. Specifically, we initialize a small test suite for a program. Then, Daikon is invoked based on the execution of this test suite to generate specifications. If the verifier cannot verify the generated specifications, we adopt a counterexample-guided strategy to include the counterexample produced by the verifier in the test suite. The procedure is repeated continuously until no new specifications are generated. As calculated by JaCoCo [41], the test suites achieve an average instruction coverage of 90.16%, branch coverage of 87.98%, and line coverage of 91.36%. Lastly, we instrument the dummy function calls at the top of each loop body in a program to ensure Houdini and Daikon can generate the specifications at these program points.
@@ IV-D Evaluation Metrics
User Rating. It aims to measure the semantic quality of the generated specifications. We invited 15 Ph.D. students who are experts in Java programming language to rate the specifications generated by different approaches. The research is conducted using the Likert Scale [48], where students are required to give a rating from one point to five points for each case. A reference rating criteria is also presented to the students.
@@ V-A RQ1: Comparison with Baselines
The experimental results are presented in Table II where Num. denotes the number of successfully handled programs, and Prob. denotes the average success probability. For the LLM-based approaches, the number of trials for each program is 10 times.
@@
Presented in Fig. 5 is the Venn diagram of the programs in SpecGenBench for which SpecGen and baseline approaches successfully generated verifiable specifications. It is noteworthy that SpecGen generates verifiable specifications for 7 programs that other baselines fail to yield, where 5 are from the Nested Loop, with the rest 2 from the Single-path Loop. Further investigation of these programs reveals that they are relatively complicated, making it challenging for the baseline approaches to be correctly generated.
@@
RQ1: SpecGen outperforms all the baselines on two different datasets, generating verifiable specifications for 279 out of 385 programs, with the highest success probability among all LLM-based approaches. In comparison, the number of programs handled by Daikon, Houdini, Conversational generation, and AutoSpec is 72, 98, 218, and 247, respectively.
@@ V-D RQ4: User Study on the Quality of Specifications
A user study is conducted to evaluate the semantic quality of the generated specifications. 15 Ph.D. students are invited to rate the specifications generated by different approaches. A detailed description of the questionnaire and rating criteria is given in Section IV-D. We selected the 15 programs from the dataset of SpecGenBench that can be handled by all of Houdini, Daikon, and SpecGen. Apart from the specifications generated by these approaches, we also add the ground truth as a reference. The specifications are kept anonymous to the students, disclosing no information about the sources of the specifications. The rating scores are presented in Table V, where the average score for full marks is 5.
@@
RQ4: SpecGen received an average rating score of 4.54, which is close to the 4.83 of the oracle specifications, demonstrating the ability of SpecGen to accurately characterize the real behaviors of the input program and generate specifications with comprehensive program semantics.
@@ VIII Conclusion
In this paper, we introduced SpecGen, a novel approach that utilizes the Large Language Model for formal program specification generation. Leveraging the code comprehension ability of LLMs as well as the well-designed mutation-based specification generation component, our approach is capable of accurately capturing the behaviour and functionality of input programs to generate accurate specifications. A comprehensive evaluation between SpecGen and other baselines is conducted on two different datasets, the benchmark for the Java category of SV-COMP, and a more diverse and manually constructed dataset containing 120 programs. The extensive experimental results haver demonstrated that our approach significantly outperforms the baseline approaches.

### v5

@@ Abstract
Index Terms:
program verification, specification inference, large language model
@@ I Introduction
  • A novel approach for formal program specification generation and corresponding prototype tool [22], leveraging the Large Language Models to generate accurate and comprehensive specifications to describe program behaviors. Benefiting from the code comprehension ability of LLMs, our approach is capable of generating specifications with high quality, overcoming the limitations of existing methods in generating simplistic and basic specifications.
@@
  • A comprehensive evaluation to evaluate our approach in all aspects. We compare SpecGen against Purely LLM-based approaches and representative non-LLM approaches. SpecGen succeeds in 279 out of the 385 programs, significantly outperforming the baseline approaches.
@@ II-A Specification Generation and Verification
Program specifications encompass precise statements that describe the intended or actual behaviors of a particular program, either in its entirety or in distinct parts. In this work, we focus on generating specifications for the actual behaviors of input programs. A large proportion of program specifications are expressed in formal languages, such as mathematical expressions to describe the constraints on the behaviors of a program. There are different kinds of specifications such as pre-conditions, which establish constraints on function parameters, ensuring proper execution of the function, post-conditions, which delineate the properties of a set of variables that persist after a function is executed, and loop invariants, which represent a specialized form of specification, detailing properties that consistently hold before executing the loop body. For different programming languages, the specifications may have different implementation forms. For example, in Java, the specifications can be expressed in Java Modeling Language (i.e., JML) [23] where requires statements denote the pre-conditions of a function, ensures statements represent the post-conditions of a function, and maintaining statements specify the loop invariants.
@@ II-B Motivation
Recently, large language models (i.e., LLMs) [28, 29] have exhibited powerful capacities in coding [19, 18]. The emergence of these models may greatly compensate for the limitations of traditional software analysis tools in code understanding. A significant amount of work attempts to leverage large language models in software engineering [30, 31, 32, 33, 34] and we have witnessed substantial progress brought about by the introduction of LLMs. Inspired by these works, in this work, we aim to leverage the large language models in the automated generation of formal program specifications to address the limitations of conventional template-based approaches. From this perspective, we innovate our approach SpecGen, which generated three parts of specifications for the example presented in the right part of Figure 1. The first part is to describe the method TwoSum() as a whole, specifying its pre-conditions and post-conditions. The specifications in lines a and b claim that the input array must not be null before and after the method is executed. The post-condition in line c specifies that the target value equals the sum of the two elements corresponding to the indexes stored in the returned array. The post-condition at line d specifies that there does not exist such a pair of elements that satisfies the constraint when the length of the returned array is zero. These generated specifications can fully articulate the functionality of method TwoSum(). Furthermore, the loop invariants in the second and third parts specify corresponding constraints that must be met within a certain range in the array. These specifications generated by SpecGen comprehensively describe the semantics of this function and their correctness is verifiable.
@@ III-A Overview
The overview of SpecGen is presented in Figure 3, which consists of two components i.e., conversation-driven specification generation and mutation-based specification generation. The former is designed to communicate with the large language model to query the output in a conversational manner. In particular, a prompt is constructed with some few-shot examples for the initial query. The verification failure information provided by the verifier is further used as the prompt for the next round of conversation if the model-generated results are incorrect. The conversation will be repeated iteratively until the generated specifications successfully pass the verifier or a maximum number of iterations is reached. The latter aims at generating the specifications of a program that the large language model fails to generate. Four kinds of mutation operators are adopted to mutate the specification that failed verification by the verifier and obtain all potential variants. A heuristic selector is further designed to efficiently choose a set of mutated variants most likely to pass verification.
@@ III-C Mutation-based Specification Generation
In the conversation-driven generation process, some few-shot examples are provided in the initial prompt to start the query with the LLM. To further stimulate the potential of the LLM, multi-turn conversation continually guides the LLM in approaching the accurate specification more closely. Yet, they still struggle to generate fully correct specifications for some complex programs. The reasons are two-fold. On the one hand, in comparison to code generation [37, 38, 39], specification generation poses greater challenges for LLMs due to the limited corpora related to program specifications for the model to learn from. On the other hand, although the verification failure information provided by the verifier can assist LLMs in providing higher-quality responses to some extent, as the error messages are highly abstract and generalized, LLMs still struggle to accurately understand the semantic information within error messages for complex programs. While LLMs may not accurately generate specifications for complex programs, the generated results are already highly close to the oracle, inspiring us to design mutation-based generation.
@@ III-C1 Template Specification Mutation
As shown in Table I, we define four kinds of mutation operators including predicative, logical, comparative, and arithmetic. Each type of mutation corresponds to one type of operator supported by JML. LLMs perform well in formulating the overall syntactical structure of specifications, but they often make mistakes in grasping the fine-grained relationships between variables, resulting in incorrect operators used to describe the relationships between variables, which is why our mutation design is centered around the operators. A mutation operation substitutes the operators of the corresponding type in the specification with another of the same type. For example, after applying a predicative mutation, a $\exists$ predicate within a specification may be substituted with $\forall$. Note that the mutation for a certain type of operator does not necessarily create only one mutated candidate. For example, the expression a <= b may be mutated to a < b or a - 1 <= b. If multiple mutations can be applied to a specification at the same time, we try to exhaust each combination of different types of mutations to get all potential variants. Since the set of all potential variants of a certain template is determined, the exhaustive searching process is deterministic. For instance, the expression x < n + 1 can be mutated to x <= n + 1 from the comparative type, x < n - 1 from the arithmetic type, or x <= n - 1 by combining them.
@@ III-C2 Mutated Specification Selection
Typically, for a program, Houdini [7] verifies all generated specifications at one time. However, similar practice cannot be applied in SpecGen as we exhaust all potential combinations of mutations for a template specification. The set of the obtained specifications for verification is considerably large, posing a much greater burden for the verifier within a single verification process. To address this challenge, we innovate a heuristic selection strategy to improve the stability and efficiency of verification.
@@ IV-B Dataset
As Daikon [9] requires a set of test suites instrumented into the source code to execute the code, we manually write these test suites for it. A small test suite is initialized first, on which Daikon is invoked to generate specifications. If the results fail verification, the counterexample produced by the verifier will be added to the test suite. The procedure is repeated until no new specifications are generated. The test suites achieve an average instruction coverage of 90.16%, branch coverage of 87.98%, and line coverage of 91.36%. Lastly, we instrument dummy function calls at the top of each loop body in a program to ensure Houdini and Daikon can generate the specifications at these program points.
@@ IV-D Evaluation Metrics
User Rating. It aims to measure the semantic quality of the generated specifications. We invited 15 Ph.D. students who are experts in Java programming language to rate the specifications generated by different approaches. The research is conducted using the Likert Scale [49], where students are required to give a rating from one point to five points for each case according to a reference rating criteria.
@@ V-A RQ1: Comparison with Baselines
The experimental results are presented in Table II where Num. denotes the number of successfully handled programs, and Prob. denotes the average success probability. Each LLM-based approach is granted 10 trials for each program.
@@
Presented in Fig. 5 is the Venn diagram of the programs in SpecGenBench for which SpecGen and baseline approaches successfully generated verifiable specifications. It is noteworthy that SpecGen generates verifiable specifications for 7 programs that other baselines fail to yield, where 5 are from the Nested Loop, with the rest 2 from the Single-path Loop. Further investigation of these programs reveals that they are relatively complicated and challenging to handle.
@@
RQ1: SpecGen outperforms all baselines on two different datasets, generating verifiable specifications for 279 out of 385 programs with the highest success probability among all LLM-based approaches. In comparison, the number of programs handled by Daikon, Houdini, Conversational generation, and AutoSpec is 72, 98, 218, and 247, respectively.
@@ V-D RQ4: User Study on the Quality of Specifications
A user study is conducted to evaluate the semantic quality of the generated specifications. 15 Ph.D. students are invited to rate the specifications generated by different approaches. A detailed description of the rating process is given in Section IV-D. We selected the 15 programs from the dataset of SpecGenBench that can be handled by all of Houdini, Daikon, and SpecGen. Apart from the specifications generated by these approaches, we also add the ground truth as a reference. The specifications are kept anonymous to the students, disclosing no information about the sources of the specifications. The rating scores are presented in Table V, where the score for full marks is 5.
@@
RQ4: SpecGen received an average rating score of 4.54, which is close to the 4.83 of the oracle specifications, demonstrating the ability to accurately characterize the real program behaviors and generate specifications with comprehensive program semantics.
@@ VIII Conclusion
In this paper, we introduced SpecGen, a novel approach that utilizes the Large Language Model for formal program specification generation. Leveraging the code comprehension ability of LLMs as well as the well-designed mutation-based specification generation component, our approach is capable of accurately capturing the behaviour and functionality of input programs to generate accurate specifications. A comprehensive evaluation between SpecGen and other baselines is conducted on two different datasets, the benchmark for the Java category of SV-COMP, and a more diverse and manually constructed dataset containing 120 programs. The extensive experimental results haver demonstrated that our approach significantly outperforms the baseline approaches, with the ability to effectively articulate program behaviors.